OpenStudy (anonymous):

use the distributive property to find each product 9*499

OpenStudy (dumbcow):

no distributive property needed here 9*499 = 4491

OpenStudy (dumbcow):

distributive property is used when you multiply a num by a sum -> 9*(x+2) = 9x +9(2) = 9x +18